A traffic helicopter pilot finds that with a tailwind her 120 km trip away from the airport takes 30 min. On her return trip to the airport, into the wind, she finds that her trip is 10 min longer. What is the speed of the helicopter? What is the speed of the wind?
Answer by mananth(16949)

let the speed of the helicopter in still air be x
let speed of wind be y
..
speed with wind = x+y
speed against wind = x-y
..
distance = speed * time
120= 1/2(x+y)
1. 240 = x+y
..
120 = 2/3*(x-y)
360 = 2x-2y
2. 180=x-y
..
Add eq1 and eq2
420=2x
x=210 kmph